My daughter is having trouble with real exponents and exponential functions in Algebra 2 could you please show me how to work this problem out. There are not enough textbooks in her class for each student to have one to bring home so therefore I have nothing to go by. Thanks.
The problem is: {(n sqrt(6) + w sqrt (3) )}x{(n sqrt (6) - w sqrt (3) )}

Found 2 solutions by Fombitz, mananth:
Answer by Fombitz(32388)

Use the FOIL method,



Answer by mananth(16949)

{(n sqrt(6) + w sqrt (3) )}x{(n sqrt (6) - w sqrt (3) )}
There is a formula for difference of squares
which is (a+b)(a-b ) = a^2-b^2
..
in this problem
a= nsqrt6
b= wsqrt3
a^2-b^2
= (n sqrt(6))^2-(wsqrt(3)^2)
= n^2*6-w^2*3
=6n^2-3w^2
=3(2n^2-w^2)
OR
...
(n sqrt(6){(n sqrt (6) - w sqrt (3)+( w sqrt (3) )}{(n sqrt (6) - w sqrt (3)
(nsqrt(6)^2-nsqrt(6)*wsqrt(3)+nsqrt(6)*wsqrt(3)-(wsqrt(3))^2
(nsqrt(6)^2-(wsqrt(3))^2
which is same as above when simplified
..
m.ananth@hotmail.ca
...
...